a boat travels 12 miles upstream and 12 miles back. the time required for the trip is 2 hours. the speed of the stream is 3 miles per hour. find the speed of the boat in still water?
Answer by nyc_function(2741)

Upstream Trip:


rate = x + 3


time = 2


distance = 12


Downstream Trip:


rate = x - 3


time = 2


distance = 12


Let x = speed of boat in still water


2(x + 3) + 2(x - 3) = 12


2x + 6 + 2x - 6 = 12


4x = 12


x = 12/4


x = 3 miles per hour